OpenStudy (anonymous):

The head circumference C of a child is related to the height H of the child (both in inches) through the function H(C) = 2.15C – 10.53 (a) Express the head circumference C as a function of height H. (b) Verify that C = C(H) is the inverse of H = H(C) by showing that H(C(H)) = H and C(H(C)) = C. (c) Predict the head circumference of a child who is 26 inches tall.

OpenStudy (anonymous):

H=2.15C-10.53 2.15C=H+10.53 C=(H+10.53)/2.15

OpenStudy (anonymous):

H(C(H)) =H((H+10.53)/2.15) =2.15(H+10.53)/2.15-10.53 =H therefore, H(C(H)) = H also, C(H(C)) =C(2.15C – 10.53) =((2.15C-10.53)+10.53)/2.15 =C therefore, C(H(C)) = C

OpenStudy (anonymous):

FROM Q1, C(H) =(H+10.53)/2.15 WHEN , H =26, C(26) =(26+10.53)/2.15 =17.0

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome!